OpenStudy (anonymous):

PLEASE HELP!Reward 1 Fan And Medal! If g(x) = x2 + 2, find g(3).

OpenStudy (catlover5925):

G(3) is 8...you plug in 3 to x then solve

OpenStudy (anonymous):

8 simply , g(3) = 2(3) + 2 = 8

OpenStudy (anonymous):

The answer is 11

OpenStudy (anonymous):

The answer is 11. You plug in 3 for the value of x. g(3) = (3)2 + 2 = 9 + 2 = 11

OpenStudy (anonymous):

how did you get 11? @MIS08

OpenStudy (anonymous):

Thank You

OpenStudy (xapproachesinfinity):

\(\huge \rm g(x)=x^2+2~~ ~~~~~~~g(\color{red}{3})=\color{red}3^2+2\)

OpenStudy (catlover5925):

its 8 if you plug in 3 to X it is 3(2) + 2 so you multiply 3 and 2 which is 6 then add 2 and you have 8

OpenStudy (anonymous):

i have one more question if you dont mind Michelle rents a movie for a flat fee of $1.50 plus an additional $1.25 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Michelle has the movie.

OpenStudy (xapproachesinfinity):

\(\huge \rm g(\square)=\square^2+2\) wherever you see X replace with what you have

OpenStudy (anonymous):

i already answered the first question i m going to ask another in anothr collum

OpenStudy (xapproachesinfinity):

another question post it in a different thread!

OpenStudy (xapproachesinfinity):

the only person who is right for the first is @MIS08 though he/she didn't provide the steps, which is not allowed here in OP